A medieval castle occupant takes a dislike to some armed visitors who are at the foot of the castle wall and drops a rock towards them if the wall is 19m high and the receivers only stands 1.6m tall how much time do they have to move aside before the rock hits one of their heads? How fast will the rock be moving if a succesful impact is made?
1
Expert's answer
2019-10-21T08:56:05-0400

The time will be

"t=\sqrt{\frac{2(H-h)}{g}}=\sqrt{\frac{2(19-1.6)}{9.8}}=1.9\ s"

The speed:


"v=\sqrt{2g(H-h)}"

"v=\sqrt{2(9.8)(19-1.6)}=18.5\frac{m}{s}"